Circular No. 7330

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gj IN NGC 3251 G. M. Hurst, Basingstoke, England, reports the discovery by Mark

Armstrong, Rolvenden, of a supernova located at α = 10^{h}29^{m}19^{s}.27, δ = +26^{o}06′19′′.2 (equinox 2000.0; average of two measurements, from Nov. 17 and Dec. 8 frames), which was noted to be 40′′ east and slightly south of the nucleus of NGC 3251. The supernova appears on three of Armstrong’s CCD frames, with the following magnitudes being estimated: Nov. 17.20 UT, 17.5; Dec. 2.14, 18.0; 8.16, 18.0. The object is not present on the first or second Palomar Sky Surveys (limiting mag ∼ 20).

P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999gj was obtained by A. Mahdavi with the F. L. Whipple Observatory (FLWO) 1.5-m Tillinghast telescope on Dec. 9.5 UT. This shows the new object to be a type-Ia event, between 1 and 2 months after maximum light. Narrow emission lines from the host galaxy give a recession velocity of 5300 km/s. CCD images obtained with the FLWO 1.2-m telescope near Dec. 9.5 yield V ∼ 19.0 for SN 1999gj. Position end figures derived from the CCD images (based on the HST Guide Star Catalog) are 19^{s}.20, 20′′.2, which is 33′′ east and 21′′ north of the galaxy nucleus.

V1494 AQUILAE M. Pontefract, Imperial College, London; R. J. Ivison, University Col-

lege, London; M. Reuland, Leiden; and R. P. J. Tilanus, Joint Astronomy Centre, Hilo, report the detection of V1494 Aql using SCUBA on the James Clerk Maxwell Telescope on Dec. 8.1 and 9.1 UT (only a week after outburst): “Measured fluxes were 19 ± 4 and 27 ± 3 mJy at 0.85 mm, with preliminary flux calibration based on Uranus. Tentative detections at 0.45 mm suggest optically thick free-free emission. The nova may soon be sufficiently bright to be detected/resolved at radio frequencies.”

Visual magnitude estimates: Dec. 5.026 UT, 4.3 (S. Giovanardi, New York City, NY); 5.643, 5.2 (K. Černis, Vilnius, Lithuania); 6.06, 5.1 (W. G. Dillon, Missouri City, TX); 6.424, 5.8 (Y.-S. Park, Chungju City, Chungbuk, South Korea); 6.742, 5.3 (S. E. Shurpakov, Baran, Belarus); 7.378, 5.6 (Y. Nagai, Yamanashi, Japan); 7.77, 5.8 (J. M. San Juan, Madrid, Spain); 8.438, 5.9 (R. Qi, Beijing, China); 8.728, 5.7 (P. Schmeer, Bischmisheim, Germany); 9.697, 5.9 (A. Baransky, Kiev, Ukraine); 10.68, 6.1 (A. Trebacz, Niepolomice, Poland); 11.403, 6.1 (S. Yoshida, Ibaraki, Japan).

1999 December 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green